# Interleaving Natural Language Prompting with Code Editing for Solving Programming Tasks with Generative AI Models

### v1

@@ 4.4. Limitations
We acknowledge several limitations of our study that could be tackled in future work. First, only the Basic Functions problems were mandatory, while more complex tasks were optional, introducing potential self-selection bias. Future studies could mitigate this by assigning tasks of varying difficulty to all students. Second, we did not conduct an A/B study to compare prompting-only and prompting-plus-editing conditions in a controlled way. A randomized study where editing capabilities are selectively enabled could more directly measure the benefits of code editing. Finally, our logs capture only edited code that was either executed or sent as an attachment, without information on how much time or effort students spent editing. Capturing intermediate drafts or more granular timestamps could provide deeper insights into students’ editing behavior.
@@ 5. Conclusion
Our study investigated a key question in AI-assisted programming: when do students revise their prompt, and when do they dive into the code? We found that students combined prompting and manual editing in their workflow, typically introducing edits after a generated solution failed execution. These manual edits were usually concise, serving as targeted, last-mile corrections. Edits became more common as tasks increased in complexity, though higher-performing students relied more consistently on prompting alone, reflecting their ability to craft effective prompts. Student reflections highlighted prompting for efficiency, editing for precision, and both as valuable tools for learning. These findings emphasize the need to support both skills in instruction, teaching students to adapt their approach dynamically. Developing this flexibility requires a solid understanding of the problem and the concepts involved in solving it.

### v2

@@ 4.4. Limitations
We acknowledge several limitations of our study that could be tackled in future work. First, only the Basic Functions problems were mandatory, while more complex tasks were optional, introducing potential self-selection bias. Future studies could mitigate this by assigning tasks of varying difficulty to all students. Second, we did not conduct an A/B study to compare prompting-only and prompting-plus-editing conditions in a controlled way. A randomized study where editing capabilities are selectively enabled could more directly measure the benefits of code editing. Third, we did not have student demographic data or their prior experience on using GenAI available to us, which limited the type of analyses we could do. Fourth, as we used a custom-built tool, it is possible the findings might not generalize to standard GenAI programming tools. Finally, our analysis spans only edited code that was either executed or sent as an attachment, excluding how much time or effort students spent editing. Analyzing intermediate drafts or more granular timestamps could provide deeper insights into students’ editing behavior.
@@ 5. Conclusion
Our study investigated a key question in AI-assisted programming: when do students revise their prompt, and when do they dive into the code? We found that students combined prompting and manual editing in their workflow, typically introducing edits after a generated solution failed execution. These manual edits were usually concise, serving as targeted, last-mile corrections. Student reflections highlighted prompting for efficiency, editing for precision, and both as valuable tools for learning. Edits became more common as tasks increased in complexity, though higher-performing students relied more consistently on prompting alone, reflecting their ability to craft effective prompts. These findings emphasize the need to support both skills in instruction, teaching students how to decide when to re-prompt versus apply an edit after testing. Tools and teaching materials can also offer scaffolding around the edit–run loop when students struggle. Developing this flexibility requires a solid understanding of the problem and the underlying programming concepts.
@@ Acknowledgements.
- Dunn (1959) Olive Jean Dunn. 1959. Estimation of the Medians for Dependent Variables. The Annals of Mathematical Statistics (1959).
@@
- Mann and Whitney (1947) Henry B Mann and Donald R Whitney. 1947. On a Test of Whether One of Two Random Variables is Stochastically Larger than the Other. The Annals of Mathematical Statistics (1947).
@@
- Tankelevitch et al. (2024) Lev Tankelevitch, Viktor Kewenig, Auste Simkute, Ava Elizabeth Scott, Advait Sarkar, Abigail Sellen, and Sean Rintel. 2024. The Metacognitive Demands and Opportunities of Generative AI. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).